# Function-Correcting Partition codes

### v1

@@ IV-A Partition graph for coset partitions
then there exists a set of representative vectors $\{u_{1},\ldots,u_{q^{\ell}}\}$ defined in the proof of Lemma 6, and $r_{\mathcal{P}}(k,t)=N(\mathcal{D}_{f}(t,u_{1},\ldots,u_{q^{\ell}}))$.
@@ IV-B Partition Graphs of Weight-Based Partitions
In this subsection, we study the partition graphs associated with structured partitions of $\mathbb{F}_{q}^{k}$ which are based on the Hamming weights of the vectors. We first consider the weight partition, denoted $\mathcal{W}=\{W_{0},W_{1},\ldots,W_{k}\}$, where each block $W_{i}$ (called the weight block) consists of all the vectors of weight $i$. We show that its partition graph contains a full-size clique, and using this property, we derive a lower bound on the redundancy of the corresponding $(\mathcal{W},t)$-FCC over $\mathbb{F}_{q}$.
We next consider a family of coarser partitions called grouped weight partitions obtained by merging weight blocks. We then show that for a subclass of it, a full size clique does not exist .
Definition 20 (Grouped weight partition).
$G_{j}=\bigcup_{i\in S_{j}}W_{i},$
A grouped weight partition $\mathcal{G}=\{G_{1},\ldots,G_{m}\}$ is called a consecutive grouped weight partition if each index set $S_{j}$ forms an integer interval, $S_{j}=\{a_{j},a_{j}+1,\ldots,b_{j}\},$ for some $0\leq a_{j}\leq b_{j}\leq k$. Equivalently, each $G_{j}$ is a union of consecutive weight blocks. The weight partition is the finest consecutive grouped weight partition.
The following lemma shows the existence of a full size clique in the partition graph of weight partition $\mathcal{W}$.
@@
Consider the weight partition $\mathcal{W}=\{W_{0},W_{1},W_{2},W_{3}\}$ of $\mathbb{F}_{3}^{3}$. Fix a nonzero symbol $a\in\mathbb{F}_{3}$, say $a=1$. Define the following vectors, as in Lemma 8:
@@
$N(\mathcal{D}_{\mathcal{W}}(t,u_{1},u_{2},u_{3},u_{4}))=\begin{bmatrix}0&4&3&2\\ 4&0&4&3\\ 3&4&0&4\\ 2&3&4&0\end{bmatrix}$.
@@
$r_{\mathcal{W}}(k,t)=N(D_{\mathcal{W}}(t\,;\,u_{0},u_{1},\ldots,u_{k})).$
Using the Plotkin bound in Lemma 1, we derive a lower bound on the redundancy of a $(\mathcal{W},t)$-encoding over $\mathbb{F}_{q}$. An upper bound is also obtained by constructing a $(\mathcal{W},t)$-encoding from a suitable error-correcting code.
@@
$\frac{2q}{(k+1)^{2}(q-1)-a(q-a)}\,S_{k,t}\leq r_{\mathcal{W}}(k,t)\leq N(\min(2t+1,k+1),2t),$
where $a=(k+1)\bmod q$, $N(M,d)$ is the minimum length of an ECC with $M$ codewords and minimum distance $d$, and
@@
$N(\mathcal{D}_{\mathcal{W}}(t;u_{0},u_{1},\ldots,u_{k}))\geq\frac{2q}{(k+1)^{2}(q-1)-a(q-a)}\sum_{i<j}\mathcal{D}_{i,j},\qquad a=(k+1)\bmod q,$
@@
Upper bound. Set $M=\min(k+1,2t+1)$, and let $\{p_{0},p_{1},\ldots,p_{M-1}\}$ be a $q$-ary error-correcting code of length $N(M,2t)$, size $M$, and minimum distance $2t$. We now construct a $\mathcal{D}_{\mathcal{W}}(t;u_{0},\ldots,u_{k})$-code of the same length.
Case 1: $k+1\leq 2t+1$ (so $M=k+1$). Define codewords $c_{i}=p_{i}$ for $i=0,1,\ldots,k$. Then for any $i\neq j$,
@@
On the other hand, by definition of the PDRM,
@@
so $d(c_{i},c_{j})\geq\mathcal{D}_{i,j}$ for all $i\neq j$. Hence $\{c_{0},\ldots,c_{k}\}$ is a $\mathcal{D}_{\mathcal{W}}(t;u_{0},\ldots,u_{k})$-code of length $N(M,2t)$.
@@
Therefore $\{c_{0},\ldots,c_{k}\}$ is a $\mathcal{D}_{\mathcal{W}}(t;u_{0},\ldots,u_{k})$-code of length $N(M,2t)$ also in this case.
Combining both cases, we get $N(\mathcal{D}_{\mathcal{W}}(t;u_{0},\ldots,u_{k}))\leq N(\min(k+1,2t+1),2t),$ and by Corollary 4 this implies
$r_{\mathcal{W}}(k,t)=N(\mathcal{D}_{\mathcal{W}}(t;u_{0},\ldots,u_{k}))\leq N(\min(k+1,2t+1),2t).$
@@
Similarly to the proof of Lemma 8, we know that $d(x,y)\geq|\mathrm{wt}(x)-\mathrm{wt}(y)|\geq|\min S_{s}-\max S_{r}|.$ For some non-zero $a\in\mathbb{F}_{q}$, we have vectors
$x=(\underbrace{a,a,\ldots,a}_{\max(S_{r})\ \text{times}},0,\ldots,0)\in W_{\max S_{r}}\subseteq G_{r}\quad\text{and}\quad y=(\underbrace{a,a,\ldots,a}_{\min(S_{s})\ \text{times}},0,\ldots,0)\in W_{\min S_{s}}\subseteq G_{s},$
@@
There are $2^{k}$ possible consecutive grouped weight partitions of $\mathbb{F}_{q}^{k}$. This is because, the weight partition $\mathcal{W}=\{W_{0},\ldots,W_{k}\}$ has $k$ potential boundaries between successive weight blocks $W_{i-1}$ and $W_{i}$, and each boundary may either be kept (separating the blocks) or removed (merging them). Thus every choice of which boundaries to keep corresponds uniquely to a consecutive grouped weight partition, giving $2^{k}$ possibilities.
In contrast, there are only $k$ distinct Hamming weight distribution functions $\Delta_{T}$ on $\mathbb{F}_{q}^{k}$, one for each threshold $T\in\{1,\ldots,k\}$. Hence Lemma 9 applies to a much larger class of weight-based partitions, and in this sense the result is strictly more general.
@@ IV-C Partition graph of the support partition
The support partition is strictly finer than the Weight partition, since each weight class $W_{i}=\{x\in\mathbb{F}_{q}^{k}:\mathrm{wt}(x)=i\}$ is subdivided into $\binom{k}{i}$ blocks $S_{A}$ of vectors having the same support. There are many functions whose domain partition is exactly $\mathcal{S}$, and an $(\mathcal{S},t)$-encoding automatically works as an $(f,t)$-FCC for all such functions. Two simple examples are:
@@
Consider space $\mathbb{F}_{3}^{3}$. For each subset $A\subseteq\{1,2,3\}$, define the block $S_{A}=\{x\in\mathbb{F}_{3}^{3}:\mathrm{supp}(x)=A\}.$ Then the support partition is $\mathcal{S}=\{S_{A}:A\subseteq\{1,2,3\}\}$ consisting of $2^{3}=8$ blocks:
@@
Now, following the construction in Lemma 10 for fix $a=1\in\mathbb{F}_{4}$, we get a clique
@@
In the following corollary, we compute a lower bound on the redundancy of a $(\mathcal{S},t)$-encoding over $\mathbb{F}_{q}$ using the Plotkin bound given in Lemma 1, and a simple upper bound using an error-correcting code.
@@
$\frac{2^{k}q}{4^{k}(q-1)-a(q-a)}\sum_{s=1}^{\min(k,2t)}(2t+1-s)\binom{k}{s}\leq r_{\mathcal{S}}(k,t)\leq N(2^{k},2t),$
where $a=2^{k}\bmod q$, and $N(M,d)$ denotes the minimum length of a $q$-ary error-correcting code with $M$ codewords and minimum distance $d$.
@@
From Corollary 6, we know that $r_{\mathcal{S}}(k,t)=N\left(\mathcal{D}_{\mathcal{S}}(t;u_{A}:A\subseteq[k])\right),$ so we need to bound $N\left(\mathcal{D}_{\mathcal{S}}(t;u_{A}:A\subseteq[k])\right)$. Let $U=\{u_{A}:A\subseteq[k]\}$ which is equivalent to space $\mathbb{F}_{2}^{k}$. If the rows and columns of the $2^{k}\times 2^{k}$ PDRM $\mathcal{D}_{\mathcal{S}}(t;u_{A}:A\subseteq[k])$ are indexed by the subsets of $[k]$, then the $(A,B)$-th entry of this PDRM is
@@
a $q$-ary error-correcting code of length $N(2^{k},2t)$, size $2^{k}$, and minimum distance $2t$ is a $\mathcal{D}_{\mathcal{S}}(t;u_{A}:A\subseteq[k])$-code. Therefore, we obtain the claimed upper bound. ∎
@@ V Block-preserving contraction of partitions
The existence of a full-size clique in the partition graph guarantees a maximal reduction of the problem of constructing an FCPC with optimal length from the entire space $\mathbb{F}_{q}^{k}$ to a set of size $E$, where $E$ denotes the number of blocks in the partition. However, such full-size cliques need not exist for an arbitrary partition, as demonstrated in Lemma 9 for certain consecutive grouped weight partitions, and even when they do, identifying them may be computationally difficult. In these situations, one may seek an intermediate solution that, while not achieving the minimal reduction to $E$, still reduces the search space to a subset of $\mathbb{F}_{q}^{k}$ that is strictly smaller than $q^{k}$ and larger than $E$.
In this section, we introduce the notion of a block-preserving contraction, which maps the entire space $\mathbb{F}_{q}^{k}$ onto a smaller subset $U$ while preserving block membership and non-increasing inter-block distances. This notion naturally generalizes the full-size clique condition, that means a full-size clique corresponds precisely to a block-preserving contraction of minimum possible size $E$. We show that whenever a block-preserving contraction exists for a partition $\mathcal{P}$, the optimal redundancy $r_{\mathcal{P}}(k,t)$ can be determined solely from the vectors in $U$ together with the induced distance constraints.
@@
Define the set $U=\{0001,\,0011,\,0111,\,1111\},$ and define $\phi:\mathbb{F}_{2}^{4}\to U$ by
@@
The following theorem shows that a block-preserving contraction is sufficient to determine the optimal redundancy.
@@
Let $N\big(\mathcal{D}_{\mathcal{P}}(t,u:u\in U)\big)=r^{\prime}$. From Theorem 2, we have $r^{\prime}\leq r_{\mathcal{P}}(k,t).$ Now we construct an $(\mathcal{P},t)$-encoding with redundancy $r^{\prime}$ to prove the reverse inequality. By the definition of $N\big(\mathcal{D}_{\mathcal{P}}(t,u:u\in U)\big)$, there exists a $\mathcal{D}_{\mathcal{P}}(t,u:u\in U)$-code $\{z_{u}:u\in U\}$ with redundancy $r^{\prime}$. We use the block-preserving contraction $(U,\phi)$ to obtain an encoding $\mathcal{C}:\mathbb{F}_{q}^{k}\rightarrow\mathbb{F}_{q}^{k+r^{\prime}}$ defined as
@@
We claim that $\mathcal{C}$ is an $(\mathcal{P},t)$-encoding. Let $x,y\in\mathbb{F}_{q}^{k}$ belong to different blocks of $\mathcal{P}$, say $x\in P_{i}$, $y\in P_{j}$ with $i\neq j$. Since $(U,\phi)$ is a block-preserving contraction, we have $\phi(x)\in P_{i},\ \phi(y)\in P_{j},$ and
@@
This holds for all $x,y$ in different blocks of $\mathcal{P}$, so $\mathcal{C}$ is an $(\mathcal{P},t)$-encoding with redundancy $r^{\prime}$. Thus
@@
In Example 12, it is not possible to reduce the problem to $E=3$ vectors corresponding to a full-size clique in the partition graph. However, the block-preserving contraction $(U,\phi)$ still allows a reduction of the problem from the entire space of $16$ vectors to a set of $4$ vectors, which serves as the next best alternative when a full-size clique does not exist.
@@
The notion of a block-preserving contraction generalizes the full-size clique condition of Theorem 7 (equivalently, the set of representatives condition for FCCs in Corollary 2). In particular, if a partition admits a block-preserving contraction $(U,\phi)$ with $|U|=E$, then the set $U$ forms a full-size clique. Other way, any full-size clique $\mathcal{C}=\{u_{1},u_{2},\ldots,u_{E}\}$ is a block-preserving contraction of minimum possible size, with the associated map $\phi_{\mathcal{C}}$ defined by $\phi_{\mathcal{C}}(x)=u_{i}$ whenever $x$ lies in the same block of the partition as $u_{i}$.
@@
Further, Lemma 11 immediately leads to the following general statements.
@@
    Let $\mathcal{G}=\{G_{1},G_{2},\ldots,G_{m}\}$ be a grouped weight partition of $\mathbb{F}_{q}^{k}$. Since the weight partition is finer than any grouped weight partition, $(U^{\prime},\phi_{U^{\prime}})$ is a block-preserving contraction of $\mathcal{G}$, where Define for $i=0,1,\ldots,k$,
@@
$r_{\mathcal{G}}(k,t)=N(\mathcal{D}_{\mathcal{G}}(t;u_{0},u_{1},\ldots,u_{k})),$
@@
Further, we give a condition for a coset partition so that it has a block-preserving contraction. For any subset $J\subseteq[k]$, define
@@
Let the coset partition of $V$ is denoted by $\mathcal{V}$.
@@
Then $\dim(V)=3$, hence the coset partition $\mathcal{V}$ of $V$ has $4$ cosets. For instance, taking representatives $00000$, $10000$, $01000$, $00100$, the four cosets are

### v2

@@ IV-A Partition graph for coset partitions
then there exists a set of representative vectors $\{u_{1},\ldots,u_{q^{\ell}}\}$ defined in the proof of Lemma 6, and $r_{\mathcal{P}}(k,t)=N_{q}(\mathcal{D}_{f}(t,u_{1},\ldots,u_{q^{\ell}}))$.
@@ IV-B Partition Graphs of Weight-Based Partitions
In this subsection, we study the partition graphs associated with partitions of $\mathbb{F}_{q}^{k}$ based on the Hamming weights of the vectors. We first consider the weight partition, $\mathcal{W}=\{W_{0},W_{1},\ldots,W_{k}\}$. We show that its partition graph contains a full-size clique, and using this property, we derive a lower bound on the redundancy of the corresponding $(\mathcal{W},t)$-FCC over $\mathbb{F}_{q}$. We next consider grouped weight partitions and show that for a subclass of it, a full size clique does not exist.
The following lemma shows the existence of a full size clique in the partition graph of the weight partition $\mathcal{W}$.
@@
Consider the weight partition $\mathcal{W}=\{W_{0},W_{1},W_{2},W_{3}\}$ of $\mathbb{F}_{3}^{3}$. Fix a nonzero symbol $a\in\mathbb{F}_{3}$, say $a=1$. Define the following vectors, as in the proof of Lemma 8:
@@
$N_{3}(\mathcal{D}_{\mathcal{W}}(t,u_{1},u_{2},u_{3},u_{4}))=\begin{bmatrix}0&4&3&2\\ 4&0&4&3\\ 3&4&0&4\\ 2&3&4&0\end{bmatrix}$.
@@
$r_{\mathcal{W}}(k,t)=N_{q}(D_{\mathcal{W}}(t\,;\,u_{0},u_{1},\ldots,u_{k})).$
Using the Plotkin bound in Lemma 1, we derive a lower bound on the redundancy of a $(\mathcal{W},t)$-encoding over $\mathbb{F}_{q}$. An upper bound is also obtained by constructing a $(\mathcal{W},t)$-encoding from a suitable ECC.
@@
$\frac{2q}{(k+1)^{2}(q-1)-a(q-a)}\,S_{k,t}\leq r_{\mathcal{W}}(k,t)\leq N_{q}(\min(2t+1,k+1),2t),$
where $a=(k+1)\bmod q$, $N_{q}(M,d)$ is the minimum length of an ECC with $M$ codewords and minimum distance $d$, and
@@
$N_{q}(\mathcal{D}_{\mathcal{W}}(t;u_{0},u_{1},\ldots,u_{k}))\geq\frac{2q}{(k+1)^{2}(q-1)-a(q-a)}\sum_{i<j}\mathcal{D}_{i,j},\qquad a=(k+1)\bmod q,$
@@
Upper bound. Set $M=\min(k+1,2t+1)$, and let $\{p_{0},p_{1},\ldots,p_{M-1}\}$ be a $q$-ary ECC of length $N_{q}(M,2t)$, size $M$, and minimum distance $2t$. We now construct a $\mathcal{D}_{\mathcal{W}}(t;u_{0},\ldots,u_{k})$-code of the same length.
Case 1: $k+1\leq 2t+1$ (so $M=k+1$). Define the codewords $c_{i}=p_{i}$ for $i=0,1,\ldots,k$. Then for any $i\neq j$,
@@
On the other hand, by the definition of the PDRM,
@@
so $d(c_{i},c_{j})\geq\mathcal{D}_{i,j}$ for all $i\neq j$. Hence $\{c_{0},\ldots,c_{k}\}$ is a $\mathcal{D}_{\mathcal{W}}(t;u_{0},\ldots,u_{k})$-code of length $N_{q}(M,2t)$.
@@
Therefore, $\{c_{0},\ldots,c_{k}\}$ is a $\mathcal{D}_{\mathcal{W}}(t;u_{0},\ldots,u_{k})$-code of length $N_{q}(M,2t)$ also in this case.
Combining both cases, we get $N_{q}(\mathcal{D}_{\mathcal{W}}(t;u_{0},\ldots,u_{k}))\leq N_{q}(\min(k+1,2t+1),2t),$ and by Corollary 5 this implies
$r_{\mathcal{W}}(k,t)=N_{q}(\mathcal{D}_{\mathcal{W}}(t;u_{0},\ldots,u_{k}))\leq N_{q}(\min(k+1,2t+1),2t).$
@@
Similarly to the proof of Lemma 8, we know that $d(x,y)\geq\mathrm{wt}(y)-\mathrm{wt}(x)\geq\min S_{s}-\max S_{r}.$ For some non-zero $a\in\mathbb{F}_{q}$, we have vectors
$x=(\underbrace{a,a,\ldots,a}_{\max S_{r}\ \text{times}},0,\ldots,0)\in W_{\max S_{r}}\subseteq G_{r}\quad\text{and}\quad y=(\underbrace{a,a,\ldots,a}_{\min S_{s}\ \text{times}},0,\ldots,0)\in W_{\min S_{s}}\subseteq G_{s},$
@@
There are $2^{k}$ possible consecutive grouped weight partitions of $\mathbb{F}_{q}^{k}$. This is because, the weight partition $\mathcal{W}=\{W_{0},\ldots,W_{k}\}$ has $k$ potential boundaries between successive weight blocks $W_{i-1}$ and $W_{i}$, and each boundary may either be kept (separating the blocks) or removed (merging them). Thus, every choice of which boundaries to keep corresponds uniquely to a consecutive grouped weight partition, giving $2^{k}$ possibilities.
In contrast, there are only $k$ distinct Hamming weight distribution functions $\Delta_{T}$ on $\mathbb{F}_{q}^{k}$, one for each threshold $T\in\{1,\ldots,k\}$. Hence, Lemma 9 applies to a much larger class of weight-based partitions, and in this sense Lemma 9 is strictly more general.
@@ IV-C Partition graph of the support partition
The support partition is strictly finer than the weight partition, since each weight class $W_{i}=\{x\in\mathbb{F}_{q}^{k}:\mathrm{wt}(x)=i\}$ is subdivided into $\binom{k}{i}$ blocks $S_{A}$ of vectors having the same support. There are many functions whose domain partition is exactly $\mathcal{S}$, and an $(\mathcal{S},t)$-encoding automatically works as an $(f,t)$-FCC for all such functions. Two simple examples are given below.
@@
Consider the space $\mathbb{F}_{3}^{3}$. For each subset $A\subseteq\{1,2,3\}$, define the block $S_{A}=\{x\in\mathbb{F}_{3}^{3}:\mathrm{supp}(x)=A\}.$ Then the support partition is $\mathcal{S}=\{S_{A}:A\subseteq\{1,2,3\}\}$ consisting of $2^{3}=8$ blocks:
@@
Now, following the construction in Lemma 10 with $a=1\in\mathbb{F}_{4}$, we get a clique
@@
In the following corollary, we compute a lower bound on the redundancy of a $(\mathcal{S},t)$-encoding over $\mathbb{F}_{q}$ using the Plotkin bound given in Lemma 1, and a simple upper bound using an ECC.
@@
$\frac{2^{k}q}{4^{k}(q-1)-a(q-a)}\sum_{s=1}^{\min(k,2t)}(2t+1-s)\binom{k}{s}\leq r_{\mathcal{S}}(k,t)\leq N_{q}(2^{k},2t),$
where $a=2^{k}\bmod q$, and $N_{q}(M,d)$ denotes the minimum length of a $q$-ary ECC with $M$ codewords and minimum distance $d$.
@@
From Corollary 7, we know that $r_{\mathcal{S}}(k,t)=N\left(\mathcal{D}_{\mathcal{S}}(t;u_{A}:A\subseteq[k])\right),$ so we need to bound $N\left(\mathcal{D}_{\mathcal{S}}(t;u_{A}:A\subseteq[k])\right)$. Let $U=\{u_{A}:A\subseteq[k]\}$ which is equivalent to the space $\mathbb{F}_{2}^{k}$. If the rows and columns of the $2^{k}\times 2^{k}$ PDRM $\mathcal{D}_{\mathcal{S}}(t;u_{A}:A\subseteq[k])$ are indexed by the subsets of $[k]$, then the $(A,B)$-th entry of this PDRM is
@@
a $q$-ary ECC of length $N_{q}(2^{k},2t)$, size $2^{k}$, and minimum distance $2t$ is a $\mathcal{D}_{\mathcal{S}}(t;u_{A}:A\subseteq[k])$-code. Therefore, we obtain the claimed upper bound. ∎
@@ V Block-preserving contraction of partitions
The existence of a full-size clique in the partition graph guaranties a maximal reduction in the size of the problem of constructing an FCPC with optimal length from the entire space $\mathbb{F}_{q}^{k}$ to a set of size $E$, where $E$ denotes the number of blocks in the partition. However, such full-size cliques need not exist for an arbitrary partition, as demonstrated in Lemma 9 for certain consecutive grouped weight partitions, and even when they do, identifying them may be computationally difficult. In these situations, one may seek an intermediate solution that, while not achieving the minimal reduction to $E$, still reduces the search space to a subset of $\mathbb{F}_{q}^{k}$ that is strictly smaller than $q^{k}$ and larger than $E$.
In this section, we introduce the notion of a block-preserving contraction, which maps the entire space $\mathbb{F}_{q}^{k}$ onto a smaller subset $U$ while preserving block membership and non-increasing inter-block distances. This notion naturally generalizes the full-size clique condition, since a full-size clique corresponds precisely to a block-preserving contraction of minimum possible size $E$. We show that whenever a block-preserving contraction exists for a partition $\mathcal{P}$, the optimal redundancy $r_{\mathcal{P}}(k,t)$ can be determined solely from the vectors in $U$ together with the induced distance constraints.
@@
Define the set $U=\{0001,\,0101,\,1101,\,1111\},$ and define $\phi:\mathbb{F}_{2}^{4}\to U$ by
@@
The following theorem shows that a block-preserving contraction is sufficient to represent the optimal redundancy.
@@
Let $N\big(\mathcal{D}_{\mathcal{P}}(t,u:u\in U)\big)=r^{\prime}$. From Theorem 2, we have $r^{\prime}\leq r_{\mathcal{P}}(k,t).$ Now we construct a $(\mathcal{P},t)$-encoding with redundancy $r^{\prime}$ to prove the reverse inequality. By the definition of $N\big(\mathcal{D}_{\mathcal{P}}(t,u:u\in U)\big)$, there exists a $\mathcal{D}_{\mathcal{P}}(t,u:u\in U)$-code $\{z_{u}:u\in U\}$ with redundancy $r^{\prime}$. We use the block-preserving contraction $(U,\phi)$ to obtain an encoding $\mathcal{C}:\mathbb{F}_{q}^{k}\rightarrow\mathbb{F}_{q}^{k+r^{\prime}}$ defined as
@@
We claim that $\mathcal{C}$ is a $(\mathcal{P},t)$-encoding. Let $x,y\in\mathbb{F}_{q}^{k}$ belong to different blocks of $\mathcal{P}$, say $x\in P_{i}$, $y\in P_{j}$ with $i\neq j$. Since $(U,\phi)$ is a block-preserving contraction, we have $\phi(x)\in P_{i},\ \phi(y)\in P_{j},$ and
@@
This holds for all $x,y$ in different blocks of $\mathcal{P}$, so $\mathcal{C}$ is a $(\mathcal{P},t)$-encoding with redundancy $r^{\prime}$. Thus
@@
In Example 14, it is not possible to reduce the problem size to $E=3$ vectors corresponding to a full-size clique in the partition graph. However, the block-preserving contraction $(U,\phi)$ still allows a reduction of the problem size from the entire space of $16$ vectors to a set of $4$ vectors, which serves as the next best alternative when a full-size clique does not exist.
@@
The notion of a block-preserving contraction generalizes the full-size clique condition of Theorem 7 (equivalently, the set of representatives condition for FCCs in Corollary 2). In particular, if a partition admits a block-preserving contraction $(U,\phi)$ with $|U|=E$, then the set $U$ forms a full-size clique. In other words, any full-size clique $\mathcal{C}=\{u_{1},u_{2},\ldots,u_{E}\}$ is a block-preserving contraction of the minimum possible size, with the associated map $\phi_{\mathcal{C}}$ defined by $\phi_{\mathcal{C}}(x)=u_{i}$ whenever $x$ lies in the same block of the partition as $u_{i}$.
@@
The Lemma 11 leads to the following general statements.
@@
    Let $\mathcal{G}=\{G_{1},G_{2},\ldots,G_{m}\}$ be a grouped weight partition of $\mathbb{F}_{q}^{k}$. Since the weight partition is finer than any grouped weight partition, $(U^{\prime},\phi_{U^{\prime}})$ is a block-preserving contraction of $\mathcal{G}$, where
@@
$r_{\mathcal{G}}(k,t)=N_{q}(\mathcal{D}_{\mathcal{G}}(t;u_{0},u_{1},\ldots,u_{k})),$
@@
Now, we give a condition for a coset partition so that it has a block-preserving contraction. For any subset $J\subseteq[k]$, define
@@
Let the coset partition of $V$ be denoted by $\mathcal{V}$.
@@
Then $\dim(V)=3$, and the coset partition $\mathcal{V}$ of $V$ has $4$ cosets. For instance, taking representatives $00000$, $10000$, $01000$, $00100$, the four cosets are